Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

no other piece of clothing is repeated?

Expert replies
by bhumika.k.shah » Sun Jan 31, 2010 1:47 am
A man chooses an outfit from 3 different shirts, 2 different pairs of shoes, and 3 different pants. If he randomly selects 1 shirt, 1 pair of shoes, and 1 pair of pants each morning for 3 days, what is the probability that he wears the same pair of shoes each day, but that no other piece of clothing is repeated?

A.(1/3)^6 * (1/2)^3
B.(1/3)^6 * (1/2)
C.(1/3)^4
D.(1/3)^2*(1/2)
E.5(1/3)^2

i tried the 3C1 approach but mid way lost track of it...
how to go about with this sum...different approaches
????
Join the discussion
Source: — Problem Solving |

by bhumika.k.shah » Sun Jan 31, 2010 1:55 am
Why will the probability on the first day be 1 ?
Join the discussion

by ajith » Sun Jan 31, 2010 1:56 am
bhumika.k.shah wrote:A man chooses an outfit from 3 different shirts, 2 different pairs of shoes, and 3 different pants. If he randomly selects 1 shirt, 1 pair of shoes, and 1 pair of pants each morning for 3 days, what is the probability that he wears the same pair of shoes each day, but that no other piece of clothing is repeated?

A.(1/3)^6 * (1/2)^3
B.(1/3)^6 * (1/2)
C.(1/3)^4
D.(1/3)^2*(1/2)
E.5(1/3)^2

i tried the 3C1 approach but mid way lost track of it...
how to go about with this sum...different approaches
????
The probability that he wears same pair shoes each day =(1/2)^2 [first day he can wear either and he wears the same next day and the day after = 1*1/2*1/2]
Probability that he wears different pairs of pants every day = 2/9
Probability that he wears different shirts every day = 2/9

Now these events are independent (ie choosing a shirt doesnt depend on what shoes is he wearing or what pants is he wearing)

Hence the probability that all these events occur is a simple multiplication of these = 4/81*1/4 = 1/81 = 1/3^4

IMO, C :)
Always borrow money from a pessimist, he doesn't expect to be paid back.
Join the discussion

by ajith » Sun Jan 31, 2010 1:57 am
bhumika.k.shah wrote:Why will the probability on the first day be 1 ?
Because on first day he has the luxury to wear any of the pairs (he either has to repeat(shoes) or not repeat (pants, shirts) in the subsequent days.
Always borrow money from a pessimist, he doesn't expect to be paid back.
Join the discussion

by bhumika.k.shah » Sun Jan 31, 2010 1:58 am
ajith i dint get it !

please explain step wise with detailed explanations :(
ajith wrote:
bhumika.k.shah wrote:A man chooses an outfit from 3 different shirts, 2 different pairs of shoes, and 3 different pants. If he randomly selects 1 shirt, 1 pair of shoes, and 1 pair of pants each morning for 3 days, what is the probability that he wears the same pair of shoes each day, but that no other piece of clothing is repeated?

A.(1/3)^6 * (1/2)^3
B.(1/3)^6 * (1/2)
C.(1/3)^4
D.(1/3)^2*(1/2)
E.5(1/3)^2

i tried the 3C1 approach but mid way lost track of it...
how to go about with this sum...different approaches
????
The probability that he wears same pair shoes each day =(1/2)^2 [first day he can wear either and he wears the same next day and the day after = 1*1/2*1/2]
Probability that he wears different pairs of pants every day = 2/9
Probability that he wears different shirts every day = 2/9

Now these events are independent (ie choosing a shirt doesnt depend on what shoes is he wearing or what pants is he wearing)

Hence the probability that all these events occur is a simple multiplication of these = 4/81*1/4 = 1/81 = 1/3^4

IMO, C :)
Join the discussion

by ajith » Sun Jan 31, 2010 2:04 am
The probability that he wears same pair shoes each day =(1/2)^2

[Say he has to pairs A and B now he can wear AAA three days or BBB three days and in total he has 8 different arrangements (2*2*2) so the probability that he wears the same shoes is 2/8 (2 favourable (AAA, BBB)/ total outcomes)]


Probability that he wears different pairs of pants every day = 2/9

[Say He has 3 Pairs of pants X, Y, Z now the total no of ways of wearing the clothes = 3*3*3

Now out which XYZ, XZY, YXZ, YZX, ZXY, ZYX are favorable so the probability = 6/27 = 2/9]


Probability that he wears different shirts every day = 2/9

[ Explanation same as that of pants]


Hope it helps!
Always borrow money from a pessimist, he doesn't expect to be paid back.
Join the discussion

by bhumika.k.shah » Sun Jan 31, 2010 2:22 am
How can this be done in the formula way ...
like 3C1 and blah!
ajith wrote:The probability that he wears same pair shoes each day =(1/2)^2

[Say he has to pairs A and B now he can wear AAA three days or BBB three days and in total he has 8 different arrangements (2*2*2) so the probability that he wears the same shoes is 2/8 (2 favourable (AAA, BBB)/ total outcomes)]


Probability that he wears different pairs of pants every day = 2/9

[Say He has 3 Pairs of pants X, Y, Z now the total no of ways of wearing the clothes = 3*3*3

Now out which XYZ, XZY, YXZ, YZX, ZXY, ZYX are favorable so the probability = 6/27 = 2/9]


Probability that he wears different shirts every day = 2/9

[ Explanation same as that of pants]


Hope it helps!
Join the discussion

by thephoenix » Sun Jan 31, 2010 2:34 am
its
for first day prob is
(3c1*2c1*3c1)/(3c1*2c1*3c1)

for second day prob is
(2c1*1c1*2c1)/(3c1*2c1*3c1)....rep is not allowed and for shoes its bcoz the same pair has to be selected

for third day prob is
1*1*1/(3c1*2c1*3c1)
final prob
[(3c1*2c1*3c1)*(2c1*1c1*2c1)*1*1*1]/[(3c1*2c1*3c1)*(3c1*2c1*3c1)*(3c1*2c1*3c1)]=1/3^4
Join the discussion

by bhumika.k.shah » Sun Jan 31, 2010 2:38 am
i like ! :-)

thank you the phoenix as usual :D
thephoenix wrote:its
for first day prob is
(3c1*2c1*3c1)/(3c1*2c1*3c1)

for second day prob is
(2c1*1c1*2c1)/(3c1*2c1*3c1)....rep is not allowed and for shoes its bcoz the same pair has to be selected

for third day prob is
1*1*1/(3c1*2c1*3c1)
final prob
[(3c1*2c1*3c1)*(2c1*1c1*2c1)*1*1*1]/[(3c1*2c1*3c1)*(3c1*2c1*3c1)*(3c1*2c1*3c1)]=1/3^4
Join the discussion